RC meeting and Run Plan, Thursday Sept. 14, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


  • Hall A call-in instructions:

- Call RC first.
- Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zak) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

  • Hall Access procedure:

- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

Readiness for Beam

    • Moller target magnet: needs to be trained for ~8h before a measurement can be done. This prevents beam delivery so we need to be careful about scheduling this.
    • SBS Left rack: breaker replaced and the setup was looked over by Joe Beaufait. At an opportune time we need facilities to fix the type of breaker they are using for our two phase supply. Alex and Ellen evened out the power draw among the two phases.

Shift schedule

  • Need TO for tomorrow swing, Sunday swing

Subsystems Checklist

For each subsystem we ask that the responsible person go over the following points:

  1. The overall status of your subsystem
  2. Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
  3. Does your system being ready depend on your team only or are outside resources required?


  • SBS magnet and Bigbite magnets
    • BigBite magnet
    • work as expected
  • Corrector magnets:
    • work as expected
  • Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
    • Laser check done today by Jack. Need to monitor the target performance and do NMR measurement
  • HCal (Jiwan Poudel)
    • VME DAC GUI does not read the correct threshold; confirm with Alex about fix
  • Bigbite (Arun Tadepalli)
    • work as expected
  • Counting House Alarms (Bill Henry)
    • Bill tested the alarms and instructions for shift crew response in the wiki (needs further revision?)
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
  • Moller polarimeter (Don Jones)
  • Hodoscopes:

Preliminary run plan

  1. First, bring the beam safely into the hall (Beam recovery procedure: https://sbs.jlab.org/wiki/index.php/HOW_TOs#Beam_Recovery_Procedure)
  2. Target performance assessment by target group, TO
  3. Reference IC calibrations
  4. Take production data if the target is ready
    1. Production: 1h for 5, 10, 20, 30uA => run 30uA until target experts confirm everything is as it should be. These will be singles (not coincidence). PS5 should be set to 2 for helicity readback.
  5. Charge asymmetry: Paul will run feedback once a day until hall C comes up.

List of items in case target isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. Moller measurement depending on availability of people and results from Hall C
  4. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)


In person: Sanghwa P, Mark J, Jiwan P, Hanjie L, Arun T, Zak. R